• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Yi Qiang

    Log entry time 02:01:57 on February 08, 2009

    Entry number 259673

    keyword=Oven temperature stablized

    The PI parameters I used for the 3 laser pumping since this afternoon
    were 300C for prop band and 600s for Rest. However, after I turned off
    the FAP laser tonight, the oven temperature started to oscillate by about
    a couple of degrees since 22:00, I then started to tweak the PI
    parameters:300C->400C->500C, 600s->900s->1000s. Right after 00:00, the
    temperature had a huge drop. I realized that the prop gain may be too
    strong and reset time may be too short, so I adjusted them to 500C and
    200s, and finally the temperature stabilized at around 02:00. Hope this
    will last long...




    Figure 1